# The rigidity of eigenfunctions’ gradient estimates

### v1

@@ 2. The sharp gradient estimates for eigenfunctions
$\nabla_{\vec{n}}|\nabla u|^{2}$ $=\sum_{i=1}^{n}2\nabla_{\vec{n}}\nabla_{i}u\cdot\nabla_{i}u=\sum_{i=1}^{n}-2(\nabla_{i}\vec{n})u\cdot\nabla_{i}u,$
$=\sum_{1\leq i,j\leq n}-2\langle\nabla_{i}\vec{n},j\rangle\nabla_{j}u\cdot\nabla_{i}u=-2\operatorname{\langle}\nabla_{\nabla u}\vec{n},\nabla u\operatorname{\rangle}\leq 0.$
@@
$\begin{split}\nabla_{\vec{n}}|\nabla u|^{2}&=\sum_{i=1}^{n}2\nabla_{\vec{n}}\nabla_{i}u\cdot\nabla_{i}u=2\nabla_{\vec{n}}\nabla_{\vec{n}}u\cdot\nabla_{\vec{n}}u=2\nabla^{2}u(e_{n},e_{n})\cdot\nabla_{\vec{n}}u\\ &=2(\Delta u-\sum_{i=1}^{n-1}\nabla^{2}u(e_{i},e_{i}))\cdot\nabla_{\vec{n}}u=-2\sum_{i=1}^{n-1}\nabla^{2}u(e_{i},e_{i})\cdot\nabla_{\vec{n}}u\\ &=-2\sum_{i=1}^{n-1}(\nabla_{i}\nabla_{i}u-(\nabla_{e_{i}}e_{i})u)\cdot\nabla_{\vec{n}}u=2\sum_{i=1}^{n-1}\operatorname{\langle}\nabla_{e_{i}}e_{i},\nabla u\operatorname{\rangle}\cdot\nabla_{\vec{n}}u\\ &=2\sum_{i=1}^{n-1}\operatorname{\langle}\nabla_{e_{i}}e_{i},e_{n}\operatorname{\rangle}\cdot(\nabla_{\vec{n}}u)^{2}\leq 0,\end{split}$
@@
(2.1) $\Psi(x)=\frac{\nabla u}{1-u^{2}}\cdot\{-\frac{1}{2}\frac{(1-u^{2})^{2}}{|\nabla u|^{4}}\cdot\nabla u\cdot\nabla\Big{(}\frac{|\nabla u|^{2}}{1-u^{2}}(x)\Big{)}-2u\}.$
@@
$\nabla F=2|\nabla u|\cdot\Big{(}\frac{\nabla^{2}u\cdot e_{1}}{1-u^{2}}+\frac{|\nabla u|^{2}\cdot u\cdot e_{1}}{(1-u^{2})^{2}}\Big{)}.$
Then we have
$u_{11}(x)=\frac{(1-u^{2})\cdot\nabla F\cdot e_{1}}{2|\nabla u|}-\frac{|\nabla u|^{2}\cdot u}{(1-u^{2})}.$
@@
(2.2) $|\nabla^{2}u|^{2}\geq u_{11}^{2}=\frac{|\nabla u|^{4}\cdot u^{2}}{(1-u^{2})^{2}}+\frac{(1-u^{2})\nabla F\cdot e_{1}}{2|\nabla u|}\cdot\Big{\{}\frac{(1-u^{2})\nabla F\cdot\nabla u}{2|\nabla u|^{2}}-2\frac{|\nabla u|^{2}u}{1-u^{2}}\Big{\}}.$
From the Bochner’s formula and (2.2), we can get
$\Delta F$ $\geq-2\nabla F\cdot\nabla[\ln(1-u^{2})]+\frac{2\cdot\Big{\{}|\nabla u|^{2}(|\nabla u|^{2}-\lambda)+(1-u^{2})\cdot(|\nabla^{2}u|^{2}-\lambda|\nabla u|^{2})\Big{\}}}{(1-u^{2})^{2}}\cdot$
$\geq\frac{2|\nabla u|^{2}}{(1-u^{2})^{2}}F-\nabla F\cdot\Psi(x);$
@@
$\nabla^{2}_{v,v}F_{\operatorname{\epsilon}}(x_{0})$ $=2\lim_{t\rightarrow 0}\frac{F_{\operatorname{\epsilon}}(\gamma_{x_{0}}(tv))-F_{\operatorname{\epsilon}}(x_{0})-\nabla_{v}F_{\operatorname{\epsilon}}(x_{0})\cdot t}{t^{2}}=2\lim_{t\rightarrow 0}\frac{F_{\operatorname{\epsilon}}(\gamma_{x_{0}}(tv))-F_{\operatorname{\epsilon}}(x_{0})}{t^{2}}$
@@ 3. The local splitting induced by eigenfunctions
$\mathcal{P}(x)=\gamma_{x}\cap\mathcal{L}_{0}:M_{p}\rightarrow\mathcal{L}_{0}$ is well-defined for any $x\in M_{p}$.
@@ 4. The global splitting
$\Psi(x)$ $=\frac{\nabla u}{1-u^{2}}\cdot\{-\frac{1}{2}\frac{(1-u^{2})^{2}}{|\nabla u|^{4}}\cdot\nabla u\cdot\nabla\Big{(}\frac{|\nabla u|^{2}}{1-u^{2}}(x)\Big{)}-2u\}$
@@
$+v_{1}\cdot(\frac{2}{x_{1}}+O(x_{1}))-\frac{2}{x_{1}^{2}}v+O(1)v$
$=2+O(x_{1})+(2x_{1}-1)\cdot(\frac{2}{x_{1}}+O(x_{1}))-\frac{2}{x_{1}^{2}}\{x_{1}^{2}-x_{1}\}$
@@
$\sup_{x\in U_{r_{0}}^{-}(N^{n-1})}(F+\epsilon\cdot v)(x)\leq\max\{0,\sup_{x\in\partial U_{r_{0}}^{-}(N^{n-1})}(F+\epsilon\cdot v)(x)\}=0.$
Note $(F+\epsilon\cdot v)(q)=0$, then we obtain
$\frac{\partial}{\partial x_{1}}\big{|}_{x=q}(F+\epsilon\cdot v)(x)\geq 0.$
@@
$\frac{\partial}{\partial x_{1}}\big{|}_{x=q}F(x)\geq-\epsilon\cdot\frac{\partial}{\partial x_{1}}\big{|}_{x=q}v(x)=\epsilon>0.$

### v2

@@ 1. Introduction
For compact Riemannian manifold $(M^{n},g)$ with $Rc\geq 0$ and smooth (possibly empty) boundary . Assume $\Delta u=-\lambda u,\displaystyle\max_{x\in M^{n}}|u(x)|=1$, where $\lambda>0$; and $|\nabla\sin^{-1}u|^{2}(p)=\lambda$ holds at some point $p$, where $|u(p)|<1$. Define $\mathcal{N}(u)=\#(M^{n}-u^{-1}(0))$ as the number of connected components of $M^{n}-u^{-1}(0)$ (the number of nodal domains of $u$); then:
@@ 2. The sharp gradient estimates for eigenfunctions
$\nabla_{\vec{n}}|\nabla u|^{2}$ $=\sum_{i=1}^{n}2\nabla_{\vec{n}}\nabla_{e_{i}}u\cdot\nabla_{e_{i}}u=\sum_{i=1}^{n}-2(\nabla_{e_{i}}\vec{n})u\cdot\nabla_{e_{i}}u,$
$=\sum_{1\leq i,j\leq n}-2\langle\nabla_{e_{i}}\vec{n},e_{j}\rangle\nabla_{e_{j}}u\cdot\nabla_{e_{i}}u=-2\operatorname{\langle}\nabla_{\nabla u}\vec{n},\nabla u\operatorname{\rangle}\leq 0.$
@@
$\begin{split}\nabla_{\vec{n}}|\nabla u|^{2}&=\sum_{i=1}^{n}2\nabla_{\vec{n}}\nabla_{e_{i}}u\cdot\nabla_{e_{i}}u=2\nabla_{\vec{n}}\nabla_{\vec{n}}u\cdot\nabla_{\vec{n}}u=2\nabla^{2}u(e_{n},e_{n})\cdot\nabla_{\vec{n}}u\\ &=2(\Delta u-\sum_{i=1}^{n-1}\nabla^{2}u(e_{i},e_{i}))\cdot\nabla_{\vec{n}}u=-2\sum_{i=1}^{n-1}\nabla^{2}u(e_{i},e_{i})\cdot\nabla_{\vec{n}}u\\ &=-2\sum_{i=1}^{n-1}(\nabla_{e_{i}}\nabla_{e_{i}}u-(\nabla_{e_{i}}e_{i})u)\cdot\nabla_{\vec{n}}u=2\sum_{i=1}^{n-1}\operatorname{\langle}\nabla_{e_{i}}e_{i},\nabla u\operatorname{\rangle}\cdot\nabla_{\vec{n}}u\\ &=2\sum_{i=1}^{n-1}\operatorname{\langle}\nabla_{e_{i}}e_{i},e_{n}\operatorname{\rangle}\cdot(\nabla_{\vec{n}}u)^{2}\leq 0,\end{split}$
@@
(2.1) $\Psi(x)=\frac{\nabla u}{1-u^{2}}\Big{\{}-\frac{1}{2}\frac{(1-u^{2})^{2}}{|\nabla u|^{4}}\nabla u\cdot\nabla\Big{(}\frac{|\nabla u|^{2}}{1-u^{2}}(x)\Big{)}-2u\Big{\}}.$
@@
$\nabla F=2|\nabla u|\Big{(}\frac{(\nabla^{2}u)e_{1}}{1-u^{2}}+\frac{u|\nabla u|^{2}e_{1}}{(1-u^{2})^{2}}\Big{)}.$
Multiplying both sides by $e_{1}$, then we have
$u_{11}(x)=\frac{(1-u^{2})(\nabla F)e_{1}}{2|\nabla u|}-\frac{|\nabla u|^{2}u}{(1-u^{2})}.$
@@
(2.2) $|\nabla^{2}u|^{2}\geq u_{11}^{2}=\frac{|\nabla u|^{4}u^{2}}{(1-u^{2})^{2}}+\frac{(1-u^{2})\nabla F\cdot e_{1}}{2|\nabla u|}\Big{\{}\frac{(1-u^{2})\nabla F\cdot\nabla u}{2|\nabla u|^{2}}-2\frac{|\nabla u|^{2}u}{1-u^{2}}\Big{\}}.$
Direct computation reveals that
(2.3) $\Delta F=\Delta\frac{|\nabla u|^{2}}{1-u^{2}}=\frac{\Delta|\nabla u|^{2}}{1-u^{2}}-\frac{|\nabla u|^{2}\Delta(1-u^{2})}{(1-u^{2})^{2}}-2\nabla\frac{|\nabla u|^{2}}{1-u^{2}}\cdot\nabla\ln(1-u^{2}).$
By the Bochner’s formula, we know
(2.4) $\Delta|\nabla u|^{2}\geq 2|\nabla^{2}u|^{2}-2\lambda|\nabla u|^{2}.$
Plugging (2.4) and (2.2) into (2.3), we have
$\Delta F$ $\geq-2\nabla F\cdot\nabla[\ln(1-u^{2})]+\frac{2\Big{\{}|\nabla u|^{2}(|\nabla u|^{2}-\lambda u^{2})+(1-u^{2})(|\nabla^{2}u|^{2}-\lambda|\nabla u|^{2})\Big{\}}}{(1-u^{2})^{2}}\cdot$
(2.5) $\geq\frac{2|\nabla u|^{2}}{(1-u^{2})^{2}}F-\nabla F\cdot\Psi(x),$
@@
$\nabla^{2}_{v,v}F_{\operatorname{\epsilon}}(x_{0})$ $=2\lim_{t\rightarrow 0}\frac{F_{\operatorname{\epsilon}}(\gamma_{x_{0}}(tv))-F_{\operatorname{\epsilon}}(x_{0})-t\nabla_{v}F_{\operatorname{\epsilon}}(x_{0})}{t^{2}}=2\lim_{t\rightarrow 0}\frac{F_{\operatorname{\epsilon}}(\gamma_{x_{0}}(tv))-F_{\operatorname{\epsilon}}(x_{0})}{t^{2}}$
@@ 3. The local splitting induced by eigenfunctions
For any $x\in M_{p}$, the set $\gamma_{x}\cap\mathcal{L}_{0}$ contains a unique point, which we denote by $\mathcal{P}(x)$.
@@ 4. The global splitting
$\Psi(x)$ $=\frac{\nabla u}{1-u^{2}}\cdot\{-\frac{1}{2}\frac{(1-u^{2})^{2}}{|\nabla u|^{4}}\nabla u\cdot\nabla\Big{(}\frac{|\nabla u|^{2}}{1-u^{2}}(x)\Big{)}-2u\}$
@@
$+v_{1}(\frac{2}{x_{1}}+O(x_{1}))-\frac{2}{x_{1}^{2}}v+O(1)v$
$=2+O(x_{1})+(2x_{1}-1)(\frac{2}{x_{1}}+O(x_{1}))-\frac{2}{x_{1}^{2}}\{x_{1}^{2}-x_{1}\}$
@@
$\sup_{x\in U_{r_{0}}^{-}(N^{n-1})}(F+\epsilon v)(x)\leq\max\{0,\sup_{x\in\partial U_{r_{0}}^{-}(N^{n-1})}(F+\epsilon v)(x)\}=0.$
Note $(F+\epsilon v)(q)=0$, then we obtain
$\frac{\partial}{\partial x_{1}}\big{|}_{x=q}(F+\epsilon v)(x)\geq 0.$
@@
$\frac{\partial}{\partial x_{1}}\big{|}_{x=q}F(x)\geq-\epsilon\frac{\partial}{\partial x_{1}}\big{|}_{x=q}v(x)=\epsilon>0.$